# Identifying Predictions That Influence the Future: Detecting Performative Concept Drift in Data Streams

### v1

@@ 6 Related Work
Perdomo et al. [6] pioneered Performative Prediction, but learning in the performative setting has emerged over the years under various names. Krempl et al. [20] coined the term influential machine learning to describe scenarios that are performative. Khritankov [21] has worked on formally modelling feedback loops that induce concept drift and other researchers such as Adam et al. [22] and Taori and Hashimoto [9] describe performative prediction scenarios as learning with feedback loops. Interestingly, their works were less concerned about finding optimal or stable points during training, but rather the real-world implications, such as model trust and bias amplification, that occur in performative settings. Performative Prediction has also expanded since its inception. Examples include Brown et al. [23] and Li and Wai [24] who have done work on state-dependent performativity while Narang et al. [25] and Piliouras and Yu [26] brought performative prediction to the multi-agent setting. Lastly, the notion of performativy is also well known in other related fields such as Recommender Systems [1] where ranking recommendations naturally influences the choices made by consumers, and Adversarial Learning [27] (e.g. Spam Detection) where detectors must adapt to malicious entities who themselves are adapting to the predictions made by the detectors.

### v2

@@ 6 Related Work
[6] pioneered Performative Prediction, but learning in the performative setting has emerged over the years under various names. [20, 21] named scenarios of prediction-induced drift as Influential Machine Learning. [22] contributed to formally modelling feedback loops that induce concept drift and other researchers such as [23, 9] describe performative prediction scenarios as learning with feedback loops. Interestingly, their works were less concerned about finding optimal or stable points during training, but rather on the real-world implications of performative settings, such as model trust and bias amplification. Performative Prediction has also expanded since its inception. Examples include [24, 25] who have done work on state-dependent performativity while [26, 27] brought performative prediction to the multi-agent setting. Lastly, the notion of performativity is also well known in other related fields such as Recommender Systems [1] where ranking recommendations naturally influences the choices made by consumers, and Adversarial Learning [28] (e.g. Spam Detection) where detectors must adapt to malicious entities who themselves are adapting to the detectors’ predictions.
@@ 7 Conclusions
Acknowledgements
We would like to thank the participants of the Dagstuhl Seminar 20372 "Beyond Adaptation: Understanding Distributional Changes" held in September 2020 for the valuable discussions on the topic of drift, and specifically on influential predictions, and the types and origins of drift. In particular, we would like to thank Battista Biggio at University of Cagliari, Matthias Deliano at Leibniz Institute for Neurobiology, Barbara Hammer at Bielefeld University, Eyke Hüllermeier at LMU Munich, Vincent Lemaire at Orange Labs, Michele Sebag at Université Paris Saclay, Myra Spiliopoulou at Magdeburg University, Jerzy Stefanowski at Poznan University of Technology, Dirk Tasche at the Swiss Financial Market Supervisory Authority, Oskar J. Gstrein and Andrej J. Zwitter at Rijksuniversiteit Groningen, and Mark Nelson and Margarita Quihuis from the Peace Innovation Lab at Stanford University.